# XAMPLER: Learning to Retrieve Cross-Lingual In-Context Examples

### v1

@@ Abstract
Recent studies have shown that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving high-quality in-context examples, significantly improves in-context learning of English. However, adapting these methods to other languages, especially low-resource ones, presents challenges due to the scarcity of available cross-lingual retrievers and annotated data. In this paper, we introduce XAMPLER: Cross-Lingual Example Retrieval, a method tailored to tackle the challenge of cross-lingual in-context learning using only annotated English data. XAMPLER first trains a retriever with positive/negative English samples, which are constructed based on the predictions of the multilingual large language model for in-context learning. Then, the trained retriever is directly employed to retrieve English examples as few-shot examples for in-context learning of target languages. Experiments on the massively multilingual text classification benchmark of SIB200 with 176 languages demonstrate that XAMPLER substantially improves the in-context learning performance across languages. Our code is available at https://github.com/cisnlp/XAMPLER.
@@ 1 Introduction
Utilizing off-the-shelf retrievers has been further validated as an effective approach in multilingual settings (Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023). However, this method encounters limitations when applied to low-resource languages. Existing multilingual retrievers, e.g., (Reimers and Gurevych, 2020), cover a limited number of languages (i.e., 50+), and language-model-based retrievers (Hu et al., 2020) struggle to effectively align distant languages (Cao et al., 2020; Liu et al., 2023). Additionally, relying on off-the-shelf retrievers might lead to sub-optimal performance. Conversely, adopting task-specific fine-tuned retrievers has been demonstrated as a more effective approach (Rubin et al., 2022). Nonetheless, the availability of data for fine-tuning task-specific retrievers in low-resource languages is limited.
To tackle these challenges, we propose a simple yet effective method that relies solely on annotated English data, termed XAMPLER (Cross-Lingual Example Retrieval). Specifically, given an English query and an English candidate training example as input, we employ in-context learning with MaLA500 (Lin et al., 2024b), a massively multilingual LLM covering 534 languages, to predict the label of the query. Based on the correctness of the prediction, we classify the candidate example as either positive or negative. Then, leveraging the curated dataset, we train a retriever based on Glot500 (Imani et al., 2023), a massively multilingual language model covering 534 languages, with the objective of minimizing the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023). Finally, the trained retriever is directly applied to retrieve valuable few-shot examples in English for the given query in the target language. The retrieved English few-shot examples, along with the input query, are then fed into MaLA500 for further in-context learning. Our extensive experiments across a range of 176 languages on SIB200 show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning performance across languages.
@@ 2.1 Problem Definition
Given an input query $q_{i}$ in any language, our aim is to retrieve tailored few-shot examples from the sample pool $D$ to enhance the in-context learning of $q_{i}$. Given the scarcity of annotated data in low-resource languages, we introduce XAMPLER, namely, Cross-Lingual Example Retrieval. On one hand, we leverage in-domain English examples as the sample pool $D$, from which we retrieve cross-lingual examples in English for $q_{i}$ in any target language. On the other hand, we only consider $q_{i}$ sourced from English training data to train the task-specific retriever, and then directly apply the retriever for evaluation across different languages.
@@ 2.2 Data Construction
To train the task-specific retriever aimed at retrieving informative samples for the given query $q_{i}$, we consider contrastive learning, which requires both positive and negative examples for each query $q_{i}$. We define examples as positive when the LLM accurately predicts the ground truth of $q_{i}$ while utilizing the candidate as a one-shot example appended to $q_{i}$ for in-context learning. Conversely, examples are categorized as negative if the LLM’s prediction deviates from the ground truth.
Scoring all pairs of training examples presents a quadratic complexity in $|D|$, making it resource-intensive. Inspired by Rubin et al. (2022), we mitigate this by selecting the top $k$ similar examples as candidates. We utilize Sentence-BERT (SBERT) (Reimers and Gurevych, 2020) [^1] This paper uses the distiluse-base-multilingual-cased-v1 version for SBERT. for candidate selection. Based on our experiments detailed in Section C, we set $k=10$. The top $k$ candidates for $q_{i}$ are denoted as $D_{i}=\{d_{i,1},\cdots,d_{i,k}\}$, where each candidate $d_{i,j}$ is represented as $(x_{i,j},y_{i,j})$, with $x_{i,j}$ being the input and $y_{i,j}$ the corresponding label.
After obtaining the candidate-query pairs $\{(q_{i},d_{i,1}),\cdots,(q_{i},d_{i,k})\}$, we conduct 1-shot in-context learning with MaLA500 (Lin et al., 2024b) to predict the class of the $q_{i}$ given the candidate $d_{i,j}$, resulting in a predicted label $\hat{y_{i,j}}$. If MaLA500 correctly predicts the label of $q_{i}$ (i.e., $\hat{y_{i,j}}=y_{i}$), we consider the candidate $d_{i,j}$ as a positive sample ($d^{+}_{i,j}$); otherwise, it is treated as a negative sample ($d^{-}_{i,j}$). Finally, we divide $D_{i}$ into sets of positive and negative samples, denoted as $D_{i}^{pos}$ and $D_{i}^{neg}$, respectively.
@@ 2.3 Retriever Fine-tuning
We utilize a contrastive loss function (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023) to train the task-specific retriever, aiming to maximize the similarity between positive example pairs while minimizing the similarity between negative example pairs. We opt for Glot500 (Imani et al., 2023) with a model size of 395M as the base model for training the retriever, considering the significant costs of fine-tuning a LLM.
@@ 2.4 In-Context Learning
When employing in-context learning across languages, where $q_{i}$ can be in any language, we utilize the fine-tuned task-specific retriever to retrieve a few cross-lingual examples in English tailored to $q_{i}$. The retrieved examples are then appended to $q_{i}$ as input for MaLA500 (Lin et al., 2024b) to predict the label of $q_{i}$ through in-context learning.
@@ 3.1 Setup
We evaluate XAMPLER on a massively multilingual text classification benchmark, i.e., SIB200 (Adelani et al., 2023). SIB200 involves seven classes, namely science/technology, travel, politics, sports, health, entertainment, and geography. Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500. For a detailed language list, refer to §A.
Our evaluation framework follows the prompt template used in Lin et al. (2024b): ‘The topic of the news [sent] is [label]’, where [sentence] represents the text for classification and [label] is the ground truth. [label] is included when the sample serves as a few-shot example but is omitted when predicting the sample. We opt for English prompt templates over in-language ones due to the labor-intensive nature of crafting templates for non-English languages, especially those with limited resources. MaLA500 takes in the concatenation of few-shot examples and $q_{i}$ as input, then proceeds to estimate the probability distribution across the label set. We measure the performance with accuracy.
@@
We compare XAMPLER with three types of baselines.
  • Random Sampling. We randomly select examples from the training set $D$.
@@
  • Cross-lingual Transfer. We delve into another category of baselines called cross-lingual transfer, which also harnesses English data. In this approach, the language model is fine-tuned with English data and subsequently deployed for evaluation across target languages. Both Glot500 and MaLA500 are included, with their corresponding cross-lingual transfer baselines are denoted as Glot500 XLT and MaLA500 XLT. For Glot500, we opt for full-parameter fine-tuning. For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we only update the parameters of the LoRA modules with prompt tuning.
In addition, we consider K-Nearest Neighbors (KNN) with XAMPLER and the baselines with retrievers for comparison. Specifically, we adopt majority voting based on the labels of the retrieved examples with the given retriever.
@@
Table 1: Average macro-accuracy across all 176 languages, 117 languages in the Latin script and 59 in non-Latin scripts on SIB200 using XAMPLER and the baselines with K-Nearest Neighbors (KNN) or In-Context Learning (ICL). The provided results are all based on the 3-shot setting.
@@ 3.2 Main Results
The comparison between the baselines and XAMPLER is illustrated in Table 1. Our analysis reveals several insights based on the performance with In-Context Learning (ICL) across different methods. Notably, the random baseline exhibits the worst performance among the baselines using ICL, emphasizing the critical role of example selection for effective in-context learning. Leveraging an off-the-shelf retriever notably boosts performance. For instance, Glot500 RET outperforms the random baseline by 9.71%, and it also outperforms MaLA500 RET, showcasing the validity of selecting Glot500 as the base retriever. Leveraging English data further enhances performance, with both Glot500 XLT and MaLA500 XLT surpassing their corresponding RET baselines by 2.66% and 8.02%, respectively.
@@ 3.4 Effect of In-Context Learning
In Table 1, we compare the performance of XAMPLER with KNN against XAMPLER with ICL. Notably, XAMPLER with ICL exhibits a superior performance, outperforming KNN by 5.63%. Moreover, XAMPLER with ICL demonstrates an average improvement of 6.54% for languages in Latin and 3.83% for languages written in non-Latin scripts. This disparity may be attributed to the ability of in-context learning to effectively model queries written in the same script as the provided examples, while facing challenges in handling queries written in different scripts from the few-shot examples.
We further compare the performance of XAMPLER with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 1. Interestingly, XAMPLER with ICL exhibits inconsistent superiority over KNN, with performance variances ranging from 3% to 10%. Specifically, XAMPLER with KNN achieves its peak performance with 5 examples, whereas ICL reaches its peak with only 2 examples. Notably, in comparison to KNN’s optimal performance, recorded at 73.26% with 5 shots, XAMPLER with ICL demonstrates a notable improvement of 2.58%. These findings underscore the efficacy of applying in-context learning in effectively leveraging the retrieved examples.
Figure 1: KNN (K-Nearest Neighbors) vs. ICL (In-Context Learning) with different number of shots. X-axis: the number of retrieved examples. Y-axis: Macro-average accuracy.
@@ 4 Related Work
While off-the-shelf retrievers have shown promise, the examples they retrieve may not always represent optimal solutions for the given task, potentially resulting in sub-optimal performance. Hence, Rubin et al. (2022) delves into learning-based approaches: if an LLM finds an example useful, the retriever should be encouraged to retrieve it. This approach enables direct training of the retriever using signals derived from query and example pairs in the task of interest.
Retrieval-based in-context learning has demonstrated effectiveness in natural language understanding (Liu et al., 2022; Ye et al., 2023), reasoning (Zhang et al., 2023), knowledge-based question answering (Liu et al., 2022; Ye et al., 2023), and text generation (Rubin et al., 2022).
A study closely related to ours is conducted by Shi et al. (2022), which trains a cross-lingual exemplar retriever via distilling the LLM’s scoring function and evaluates it on four languages for the Text-to-SQL Semantic Parsing task. However, our contribution lies in addressing the more challenging low-resource scenario, thereby extending the applicability and robustness of the approach proposed by Shi et al. (2022).
@@ 5 Conclusion
In this paper, we introduce XAMPLER, a novel approach designed for cross-lingual example retrieval to facilitate in-context learning in any language. By relying solely on English data, XAMPLER trains a task-specific retriever capable of retrieving cross-lingual English examples tailored to any language query, thereby facilitating few-shot in-context learning for any language. Our experiments on SIB200 across 176 languages show that XAMPLER can outperform previous methods by a notable margin.
Limitation
In this paper, we utilize Glot500 as the base model for training XAMPLER’s retriever, MaLA500 as the LLM for in-context learning, and SIB200 as the benchmark to assess the efficacy of XAMPLER across multiple languages. We opted not to consider alternative models and benchmarks due to the absence of massively multilingual models and benchmarks.
Additionally, while it is acknowledged that English may not universally serve as the optimal source language for cross-lingual transfer across all target languages (Lin et al., 2019; Wang et al., 2023; Lin et al., 2024a), our study does not explore the selection of different source languages due to the predominant availability of training data in English for many tasks.
@@ Appendix B KNN Performance Across Layers
We show the 10-shot KNN results across layers with Glot500 and MaLA500 as retrievers in Figure 2 and 3. As shown, layer 21 of MaLA500 and layer 11 of Glot500 achieve the best performance across layers. Therefore, the retrieved results based on these two layers are used in the baselines.

### v2

@@ Abstract
Recent studies indicate that leveraging off-the-shelf or fine-tuned retrievers, capable of retrieving relevant in-context examples tailored to the input query, enhances few-shot in-context learning of English. However, adapting these methods to other languages, especially low-resource ones, poses challenges due to the scarcity of cross-lingual retrievers and annotated data. Thus, we introduce XAMPLER: Cross-Lingual Example Retrieval, a method tailored to tackle the challenge of cross-lingual in-context learning using only annotated English data. XAMPLER first trains a retriever based on Glot500, a multilingual small language model, using positive and negative English examples constructed from the predictions of a multilingual large language model, i.e., MaLA500. Leveraging the cross-lingual capacity of the retriever, it can directly retrieve English examples as few-shot examples for in-context learning of target languages. Experiments on the multilingual text classification benchmark SIB200 with 176 languages show that XAMPLER substantially improves the in-context learning performance across languages. Our code is available at https://github.com/cisnlp/XAMPLER.
XAMPLER: Learning to Retrieve Cross-Lingual In-Context Examples
Peiqin Lin1,2, André F. T. Martins3,4,5, Hinrich Schütze1,2 1Center for Information and Language Processing, LMU Munich 2Munich Center for Machine Learning 3Instituto Superior Técnico, Universidade de Lisboa (Lisbon ELLIS Unit) 4Instituto de Telecomunicações  5Unbabel linpq@cis.lmu.de
@@ 1 Introduction
Figure 1: XAMPLER involves three steps: 1. Data Construction: given a query in English $q_{i}$, we divide the candidate examples $D_{i}$ into positive examples $D_{i}^{pos}$ and negative examples $D_{i}^{neg}$ based on the prediction of MaLA500 (Lin et al., 2024b); 2. Retriever Fine-tuning: we fine-tune the retriever based on Glot500 (Imani et al., 2023) using the constructed data; 3. In-Context Learning: given a query in any language $q_{j}$, we use the fine-tuned retriever to retrieve relevant English examples as few-shots for in-context learning. During training, XAMPLER is English-only, but for evaluation via in-context learning, it extends to any of the 500+ languages supported in MaLA500 and Glot500.
Utilizing off-the-shelf retrievers has been further validated as an effective approach in multilingual settings (Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023). However, this method encounters limitations when applied to low-resource languages. Existing multilingual retrievers, e.g., SBERT (Reimers and Gurevych, 2020), cover a limited number of languages (i.e., 50+), and language-model-based retrievers (Hu et al., 2020) struggle to effectively align distant languages (Cao et al., 2020; Liu et al., 2023). Additionally, relying on off-the-shelf retrievers might lead to sub-optimal performance. Conversely, adopting task-specific fine-tuned retrievers has been demonstrated as a more effective approach (Rubin et al., 2022). Nonetheless, the availability of data for fine-tuning task-specific retrievers in low-resource languages is limited.
To tackle these challenges, we propose a simple yet effective method that relies solely on annotated English data, termed XAMPLER (Cross-Lingual Example Retrieval). As shown in Fig. 1, given an English query $q_{i}$ and an English example from the candidate pool $D_{i}$, we employ in-context learning with MaLA500 (Lin et al., 2024b), a 10B multilingual LLM covering 534 languages, to predict the label of the query. Based on the correctness of the prediction, we classify the candidate example as either positive or negative, i.e., $D_{i}^{pos}$ and $D_{i}^{neg}$. Then, leveraging the curated dataset, we train a retriever based on Glot500 (Imani et al., 2023), a multilingual small language model covering 534 languages, aiming to minimize the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023). Finally, the trained retriever is directly applied to retrieve valuable few-shot examples in English for the given query in the target language. The retrieved English few-shot examples, along with the input query, are then fed into MaLA500 for in-context learning. Experiments across 176 languages on SIB200 show that XAMPLER effectively retrieves cross-lingual examples, thereby enhancing in-context learning across languages.
@@ 2.1 Problem Definition
Given an input query $q_{i}$ in any language, our objective is to enhance in-context learning for predicting the label of $q_{i}$ by retrieving tailored few-shot examples from the pool of candidate examples $D$. Due to the scarcity of annotated data in low-resource languages, we introduce XAMPLER, namely, Cross-Lingual Example Retrieval. On one hand, we leverage in-domain English examples as the pool of candidate examples $D$, from which we retrieve cross-lingual examples in English for $q_{i}$ in any target language. On the other hand, we only consider $q_{i}$ sourced from English training data to train the task-specific retriever, which is then directly applied for evaluation across languages.
@@ 2.2 Data Construction
To train the task-specific retriever aimed at retrieving informative examples for the given query $q_{i}$, we consider contrastive learning, which requires both positive and negative examples for each query $q_{i}$. We define examples as positive when the LLM accurately predicts the ground truth of $q_{i}$ while utilizing the example as a one-shot example appended to $q_{i}$ for in-context learning. Conversely, examples are categorized as negative if the LLM’s prediction deviates from the ground truth.
Scoring all pairs of training examples presents a quadratic complexity in $|D|$, making it resource-intensive. Inspired by Rubin et al. (2022), we mitigate this by selecting the top $k$ similar examples as candidates. We utilize Sentence-BERT (SBERT) (Reimers and Gurevych, 2020) [^1] We use version distiluse-base-multilingual-cased-v1. for candidate selection. Based on our experiments detailed in Section C, we set $k=10$. The top $k$ candidates for $q_{i}$ are denoted as $D_{i}=\{d_{i,1},\cdots,d_{i,k}\}$, where each candidate $d_{i,j}$ is represented as $(x_{i,j},y_{i,j})$, with $x_{i,j}$ being the input and $y_{i,j}$ the corresponding label.
After obtaining the candidate-query pairs $\{(q_{i},d_{i,1}),\cdots,(q_{i},d_{i,k})\}$, we conduct 1-shot in-context learning with MaLA500 (Lin et al., 2024b) to predict the class of the $q_{i}$ given the candidate $d_{i,j}$, resulting in a predicted label $\hat{y}_{i,j}$. If MaLA500 correctly predicts the label of $q_{i}$ (i.e., $\hat{y}_{i,j}=y_{i}$), we consider the candidate $d_{i,j}$ as a positive example ($d^{+}_{i,j}$); otherwise a negative example ($d^{-}_{i,j}$). Finally, we divide $D_{i}$ into sets of positive and negative examples, denoted as $D_{i}^{pos}$ and $D_{i}^{neg}$, respectively.
@@ 2.3 Retriever Fine-tuning
We utilize the contrastive loss (Rubin et al., 2022; Cheng et al., 2023; Luo et al., 2023) to train the task-specific retriever, aiming to maximize the similarity between $q_{i}$ and $x_{i,j}$ if $x_{i,j}$ is a positive example while minimizing the similarity if $x_{i,j}$ is a negative example. We opt for Glot500 (Imani et al., 2023) with a model size of 395M as the base model for training the retriever, considering the significant cost of fine-tuning an LLM. We train for 50 epochs with a learning rate of 2e-5. Due to the multilingual nature of Glot500, the fine-tuned retriever can be effectively transferred to retrieve in-context examples for other languages.
@@ 2.4 In-Context Learning
At test time, when employing in-context learning across languages, where $q_{i}$ can be in any language, we use the fine-tuned task-specific retriever to retrieve a few cross-lingual examples in English tailored to $q_{i}$. The retrieved examples are appended to $q_{i}$ as input for MaLA500 (Lin et al., 2024b) to predict the label of $q_{i}$ through in-context learning.
@@ 3.1 Setup
We evaluate XAMPLER on a massively multilingual text classification benchmark, SIB200 (Adelani et al., 2023). SIB200 involves seven classes: science/technology, travel, politics, sports, health, entertainment, and geography. Our evaluation spans a diverse set of 176 languages, obtained by intersecting the language sets of SIB200 and MaLA500 (see §A). The English training set contains 701 samples, and each language has 204 samples for evaluation.
Our evaluation framework follows the prompt template used in Lin et al. (2024b): ‘The topic of the news [sent] is [label]’, where [sentence] represents the text for classification and [label] is the ground truth. [label] is included when the sample serves as a few-shot example but is omitted when predicting the sample. We opt for English prompt templates over in-language ones due to the labor-intensive nature of crafting templates for non-English languages, especially those with limited resources. MaLA500 takes the concatenation of few-shot examples and $q_{i}$ as input, then proceeds to estimate the probability distribution across the label set. We measure the performance with accuracy.
@@
Random Sampling. We randomly select examples from the English candidate pool $D$.
@@
Cross-lingual Transfer. Cross-lingual transfer is another baseline exploiting English data. In this approach, the multilingual language model is fine-tuned with English data and then deployed for evaluation across target languages. Both Glot500 and MaLA500 are included, with their corresponding cross-lingual transfer baselines denoted as Glot500 XLT and MaLA500 XLT. For Glot500, we opt for full-parameter fine-tuning. For MaLA500, which is trained by incorporating LoRA (Hu et al., 2022) into LLaMA 2-7B (Touvron et al., 2023), we only update the LoRA parameters with prompt tuning.
K-Nearest Neighbors. We consider K-Nearest Neighbors (KNN) with the fine-tuned task-specific retriever of XAMPLER and the baselines with retrievers for comparison. Specifically, we adopt majority voting based on the labels of the examples retrieved by the given retriever.
@@
Table 1: Average macro-accuracy across all 176 languages, 117 languages in Latin scripts and 59 in non-Latin scripts on SIB200 using XAMPLER and the baselines. KNN results are only provided for retriever-based methods. Results are based on the 3-shot setting.
@@ 3.2 Main Results
The comparison between the baselines and XAMPLER is illustrated in Table 1. Our analysis reveals several insights based on the performance with In-Context Learning (ICL) across different methods. Notably, the random baseline exhibits the worst performance among the baselines using ICL, emphasizing the critical role of example selection for effective in-context learning. Leveraging an off-the-shelf retriever notably boosts performance. For instance, Glot500 RET outperforms the random baseline by 9.71%, and it also outperforms MaLA500 RET, showcasing the validity of selecting Glot500 as the base retriever. Leveraging English data further enhances performance, with both Glot500 XLT and MaLA500 XLT surpassing their corresponding RET baselines by 2.66% and 8.02%.
@@ 3.4 Effect of In-Context Learning
In Table 1, we compare XAMPLER with KNN against that with ICL. Notably, XAMPLER with ICL surpasses XAMPLER with KNN by 5.63%. Moreover, XAMPLER with ICL demonstrates an average improvement of 6.54% for languages in Latin and 3.83% for languages in non-Latin. This disparity may be attributed to the ability of in-context learning to effectively model queries written in the same script as the provided examples, while facing challenges in handling queries written in different scripts from the few-shot examples.
We further compare XAMPLER’s performance with KNN and ICL using varying numbers of retrieved examples, as illustrated in Figure 2. Interestingly, XAMPLER with ICL exhibits inconsistent superiority over KNN, with performance variances ranging from 3% to 10%. Specifically, XAMPLER with KNN achieves its peak performance with 5 examples, whereas ICL achieves impressive results with only 2 examples. Notably, in comparison to KNN’s optimal performance, recorded at 73.26% with 5 shots, XAMPLER with ICL demonstrates a notable improvement of 2.58%. These findings underscore the efficacy of applying in-context learning in effectively leveraging the retrieved examples.
Figure 2: KNN (K-Nearest Neighbors) vs. ICL (In-Context Learning) with different number of shots. X-axis: number of shots. Y-axis: Macro-average accuracy.
@@ 4 Related Work
While off-the-shelf retrievers have shown promise, the examples they retrieve may not always represent optimal solutions for the given task, potentially resulting in sub-optimal performance. Hence, Rubin et al. (2022) delve into learning-based approaches: if an LLM finds an example useful, the retriever should be encouraged to retrieve it. This approach enables direct training of the retriever using signals derived from query and example pairs in the task of interest.
Several works (Shi et al., 2022; Nie et al., 2022; Winata et al., 2023; Tanwar et al., 2023) extend these methods to non-English languages. A study closely related to ours is Shi et al. (2022), which trains a cross-lingual example retriever via distilling the LLM’s scoring function and evaluates it on four languages for the Text-to-SQL Semantic Parsing task. However, our contribution lies in addressing the more challenging low-resource scenario, thereby extending the applicability and robustness of the approach proposed by Shi et al. (2022).
@@ 5 Conclusion
In this paper, we introduce XAMPLER, a novel approach designed for cross-lingual example retrieval to facilitate in-context learning in any language. Relying solely on English data, XAMPLER trains a task-specific retriever capable of retrieving cross-lingual English examples tailored to any language query, thereby facilitating few-shot in-context learning for any language. Our experiments on SIB200 across 176 languages show that XAMPLER can outperform previous methods by a notable margin.
Limitations
We did not consider other models and benchmarks due to the absence / unavailability of massively multilingual ones. Additionally, while it is acknowledged that English may not universally serve as the optimal source language for cross-lingual transfer across all target languages (Lin et al., 2019; Wang et al., 2023; Lin et al., 2024a), our study does not explore the selection of different source languages due to the predominant availability of training data in English for many tasks.
@@ Appendix B KNN Performance Across Layers
We show the 10-shot KNN results across layers with Glot500 and MaLA500 as retrievers in Figure 3 and 4. As shown, layer 21 of MaLA500 and layer 11 of Glot500 achieve the best performance across layers. Therefore, the retrieved results based on these two layers are used in the baselines.